### v1

# Forging Time Series with Language: A Large Language Model Approach to Synthetic Data Generation

Abstract
SDForger is a flexible and efficient framework for generating high-quality multivariate time series using LLMs. Leveraging a compact data representation, SDForger provides synthetic time series generation from a few samples and low-computation fine-tuning of any autoregressive LLM. Specifically, the framework transforms univariate and multivariate signals into tabular embeddings, which are then encoded into text and used to fine-tune the LLM. At inference, new textual embeddings are sampled and decoded into synthetic time series that retain the original data’s statistical properties and temporal dynamics. Across a diverse range of datasets, SDForger outperforms existing generative models in many scenarios, both in similarity-based evaluations and downstream forecasting tasks. By enabling textual conditioning in the generation process, SDForger paves the way for multimodal modeling and the streamlined integration of time series with textual information. SDForger source code will be open-sourced soon.
1 Introduction
In the era of foundation models, integrating time-series analysis with language models (LLMs) has emerged as a key research priority, spurring work on specialized models for forecasting, anomaly detection and root cause analysis (Liang et al., 2024). These efforts aim to harness the representational power of LLMs to tackle complex temporal dependencies. Despite these advancements, foundation models for time series still struggle with concept drift, high variability, and face performance degradation over long horizons. In addition, the scarcity of large and diverse time-series datasets often limits model generalization, especially in domains where data collection is expensive or operationally constrained, such as climate science, finance, or plasma physics (Kit et al., 2024).
In this context, synthetic data generation has emerged as a complementary research direction to improve scalability and model performance, particularly in low-quality or data-scarce settings. In particular, synthetic data can be leveraged to fine-tune machine learning models on domain-specific distributions, reducing the need to train models from scratch and improving sample efficiency. Indeed, several generative models for time series have been proposed, including VAE-based architectures  (Desai et al., 2021; Lee et al., 2023), GANs (Pei et al., 2021; Kidger et al., 2021), and diffusion-based approaches (Zhou et al., 2023). While these methods have shown promise, they typically lack pretraining, require task-specific retraining, and often struggle with long-term dependencies, multivariate coupling, and distributional shifts (Ang et al., 2023).
Recent studies have instead explored the use of LLMs for synthetic tabular data generation (Padhi et al., 2021; Borisov et al., 2022), demonstrating that language models trained on text-encoded data can capture statistical relationships and feature dependencies effectively. However, extending these methods to time series is non-trivial: long temporal windows increase inference and training costs, while temporal and multivariate correlations require more structured modeling.
These limitations highlight the need for novel methodologies that adapt LLMs for time-series generation while addressing temporal, structural, and computational challenges. We introduce SDForger (Synthetic Data Forger), a novel framework for generating high-quality univariate and multivariate time series, even in data-scarce settings. SDForger leverages foundation models with minimal fine-tuning by operating over compact tabular embeddings derived from functional decompositions.
Key features and advantages of SDForger:
  • Compact basis representation SDForger uses FastICA or PCA to embed time series into low-dimensional tabular data. These capture key temporal and inter-variable structures while decoupling the embedding from sequence length, enabling efficient processing of long signals.
  • Text-to-sequence generation via LLMs The embedding tables are converted into structured textual prompts and used to fine-tune a language model. A guided inference approach is then used to generate structured embeddings, ensuring that the synthetic data retains the original dataset’s statistical properties and feature relationships.
  • Flexible, lightweight architecture The framework leverages autoregressive LLMs, including lightweight models, and requires only a small number of training instances. Its modular design enables easy adaptation to different generation tasks and architectures, while its compact embedding space ensures fast inference, even for long time-series windows.
  • Multivariate and multimodal readiness SDForger can model complex multivariate dynamics and supports future extensions to textual conditioning, enabling generation guided by both time-series structure and external language-based context.
By combining structured embeddings with LLM-based generation, SDForger establishes a new paradigm for scalable, interpretable, and high-quality synthetic time-series generation.
Our simulations demonstrate that SDForger not only generates statistically realistic time series but also improves downstream model performance, often matching or exceeding results obtained from real data alone. This is particularly valuable in practical scenarios where access to high-quality data is limited or where distribution shifts make original training data less effective. Compared to state-of-the-art generative models, SDForger achieves competitive or superior performance across a wide range of similarity metrics and utility-based evaluations. Notably, our experiments show that even lightweight, pretrained LLMs (e.g., GPT-2) are sufficient to produce high-quality synthetic data with minimal fine-tuning, highlighting the accessibility, efficiency, and flexibility of our approach.
In the remainder of this paper, we first review related work (Section 2), then detail the SDForger framework (Section 3), describe our evaluation setup (Section 4), present extensive experimental results (Section 5), highlight the flexibility of language models (Section 6), and conclude with key takeaways and future directions (Section 7).
2 Related work
Time-series generation
Recent advances in time-series generation have introduced a variety of deep generative models, including GAN-based approaches like TimeGAN (Smith and Smith, 2020), state-space models, and vector-quantized architectures such as TimeVQVAE (Lee et al., 2023). While these methods generate realistic sequences, they often struggle with multivariate dependencies, and require training from scratch for each dataset. More recent framework integrate randomly-weighted combinations of time series to improve their pretraining pipeline, e.g. Chronos (Ansari et al., 2024) adapting the Mixup (Zhou et al., 2023) methodology for time series. To address evaluation challenges, TSGBench (Ang et al., 2023) proposes a unified benchmark of similarity, fidelity, and utility metrics.
Foundation models and LLMs for time series
Recent advances have introduced foundation models specifically designed for time-series tasks, offering unified frameworks for forecasting, classification, and anomaly detection (Liang et al., 2024). Examples include TFT (Lim et al., 2021), TimeGPT (Garza and Mergenthaler-Canseco, 2023), and Chronos (Ansari et al., 2024), while lightweight models like TTMs (Ekambaram et al., 2024) focus on efficient multivariate forecasting. Parallel efforts have explored adapting LLMs to time-series data by encoding numerical sequences as text (Gruver et al., 2024; Zhou et al., 2023; Jin et al., 2023), enabling zero-shot inference and transfer learning. Notably, LLMTime (Gruver et al., 2024) and GPT4TS (Zhou et al., 2023) retain most of the LLM architecture while fine-tuning only shallow layers, and Time-LLM (Jin et al., 2023) employs reprogramming to adapt to temporal tasks. Despite these innovations, most existing approaches either require task-specific pretraining or struggle to model complex structures and may benefit from synthetic data for improving generalization and robustness.
3 Methodology
In this section, we present our methodology, illustrated in Figure 1. SDForger is divided into three macro-steps: (i) Preprocessing and embedding transform the time series into tabular data (i.e., steps 1 to 3 highlighted in purple); (ii) Fine-tuning and Generation fine-tune a pre-trained LLM and generate new embedding instances (i.e., step 4 highlighted in green); (iii) Decoding reconstruct the original time-series space from the generated embeddings (i.e., steps 5 and 6 highlighted in light blue).
Figure 1: SDForger pipeline. Overview of the SDForger generation process. The example illustrates a setting with $I=3$ input segments, $C=2$ channels, $k_{1}=k_{2}=3$ components, and $\tilde{I}=2$ generated samples. The model performs periodicity-aware segmentation, extracts embeddings, and embed them into text. An LLM is then fine-tuned to generate embedding sequences, which are finally decoded to reconstruct synthetic time series.
Notation
Hereinafter, we introduce some basic notation. Let $X=\{X_{i}\}_{i=1}^{I},\ {X_{i}}\in\mathbb{R}^{C\times L}$ represent a collection of $I$ instances of a multivariate time series, where each instance has length $L$ and consists of $C$ channels. The task of synthetic time-series generation can now be formally defined as producing $\tilde{I}$ instances of a multivariate time series $\{\tilde{X}_{i}\}_{i=1}^{\tilde{I}},\tilde{X}_{i}\in\mathbb{R}^{C\times L}$ conditioned on the given context $X$. Throughout the paper, we denote by $x_{i}^{c}\in\mathbb{R}^{L}$ the $i$-th instance of channel $c$, and by $X^{c}\in\mathbb{R}^{I\times L}$ the matrix collecting all instances associated with channel $c$.
Periodicity-aware segmentation
In cases where only a single time series instance $X_{1}\in\mathbb{R}^{C\times L_{1}}$ is available, we apply a segmentation strategy to artificially create multiple instances. The segmentation procedure facilitates the estimation of the time-series distribution. Specifically, we extract $I$ periodicity-aware windows of fixed length $L<L_{1}$, aligning cuts with natural cycles and minimizing overlap to enhance independence and diversity. This pre-processing (described in Appendix A.1) transforms the data from a single sequence $X_{1}$ into a set $X=\{X_{i}\}_{i=1}^{I}$, where $X_{i}\in\mathbb{R}^{C\times L}$, preparing the data for the generation task.
3.1 From time series to tabular data
To enable tabular generation and analysis, SDForger transforms time series into structured tabular data using basis decomposition techniques. Each row represents a time series embedding obtained by projecting the signal onto a set of learned basis functions. Specifically, we adopt two decomposition methods: Functional Principal Components (FPC) (Ramsay and Silverman, 2005) and Fast Independent Component Analysis (FastICA) (Hyvarinen, 1999):
  • FPC identifies principal modes of variation by performing eigen-decomposition of the covariance operator. It captures directions of maximal variance, preserving correlation across components, showing effectiveness in modeling multivariate longitudinal data (Boschi et al., 2024).
  • FastICA extracts statistically independent components by maximizing non-Gaussianity. It decomposes a contrast function, uncovering independent latent factors that may not align with the directions of maximal variance.
Formally, for each channel $c$, we assume the instances $(X^{c}_{1},\dots,X^{c}_{I})$ are realizations of continuous functions defined over $\mathcal{T}=[0,L]$. We approximate each $X^{c}_{i}$ as a linear combination of $k_{c}$ basis functions $(b^{c}_{1},\dots,b^{c}_{k_{c}})$, where the choice of basis depends on the decomposition method (non-Gaussianity-based for FastICA, covariance-based for FPC). The embedding coefficients are:
$e^{c}_{ij}=\langle X^{c}_{i},b^{c}_{j}\rangle_{\mathbb{L}^{2}}=\int_{\mathcal{T}}X^{c}_{i}(t)b^{c}_{j}(t)\,\mathrm{d}t,$
We define the embedding matrix for channel $c$, $E^{c}\in\mathbb{R}^{I\times k_{c}}$. By concatenating the embeddings across all channels, we obtain the final embedding table: $E=(E^{1},\dots,E^{C})\in\mathbb{R}^{I\times K}$ with $K=\sum_{c=1}^{C}k_{c}.$
Throughout this paper, we refer to the columns of $E$ as embedding features. We denote the $i$-th row of $E$ as $E_{i}$, corresponding to the embedding vector of instance $X_{i}$, and the value in its $k$-th column as $e_{ik}$. More details on the choice of $k_{c}$ are given in Appendix A.2.
Notably, both methods offer the advantage that their computational cost depends on the number of instances $I$ and the number of components $k_{c}$, but not on the instance length $L$. This decoupling allows our algorithm to handle very long time windows without a corresponding increase in computational complexity, ensuring great flexibility and scalability.
3.2 Generation of tabular data
Our data generation block consists of three key stages: encoding tabular data into text, fine-tuning an LLM, and generating synthetic embeddings.
3.2.1 From embeddings table to text
LLMs are designed to process textual information. Therefore, applying an LLM to tabular data requires converting each row into a textual format that can serve as a prompt during the fine-tuning stage. Inspired by Donahue et al. (2020), we introduce a Textual Encoder responsible for converting tabular instances $E_{i}$ into structured text representations using a Fill-In-The-Middle template.
Definition 1 (Textual encoder)
Let $\mathcal{P}^{\textup{FT}}=\{{\mathcal{P}_{i}^{\textup{FT}}\}}_{i=1}^{I}$ denote the set of fine-tuning prompts, where:
[…]
Random Feature Order Permutation. Encoding tabular data into text can introduce unintended positional biases, as LLMs inherently process tokens in sequence. To enforce order independence (Borisov et al., 2022), we apply a random permutation $\pi$ to the encoded feature-value pairs within each instance. This shuffling ensures that the model does not infer any spurious relationships based on the ordering of features within the textual representation. For $K=2$, an admissible finetuning prompt for $E_{i}$ is: ‘‘Input: value_2 is [blank], value_1 is [blank] [sep] Target: $e_{i2}$ [answer] $e_{i1}$ [answer]’’
3.2.2 Large language model finetuning and inference
Fine-tuning
By training an LLM on structured text representations of the embedding tables, we enable it to learn meaningful patterns present in the data. Since the optimal number of fine-tuning epochs depends on the number of instances, the embedding dimension, and the LLM architecture, we implement an early stopping criterion to prevent overfitting.
Inference
After fine-tuning, inference is performed by prompting the LLM with structured textual templates that mirror the training format, allowing it to autonomously generate new embedding rows.
Definition 2 (Textual inference)
Given the embedding table $E\in\mathbb{R}^{I\times K}$, we define the set of inference prompts at each generation step as $\mathcal{P}^{\textup{INF}}=\{{\mathcal{P}_{g}^{\textup{INF}}\}}_{g=1}^{G}$ where:
[…]
We use a multinomial distribution sampling strategy to reduce repetition and generate more creative and diverse outputs. The model draws from its learned token probability distribution at each step, guided by the temperature parameter, which controls sampling variability. As a result, all values are internally generated by the LLM in a fully conditional and self-contained manner, highlighting the model capacity to internalize statistical and structural patterns from compact embeddings and synthesize coherent time series without external noise injection or sampling routines.
At each inference step, we generate a batch of $G$ synthetic instances, repeating the process until the desired number of sequences is obtained or a stopping criterion is met. We denote the set of all generated text instances as: $\mathcal{G}=\{\mathcal{G}_{1},\dots,\mathcal{G}_{G}\}.$ Ideally, the fine-tuned LLM should generate text instances in the following format: $\mathcal{G}_{g}=\bigcirc\ (\ \mathcal{P}_{g}^{\textup{INF}},\ \bigcirc_{k=1}^{K}(\text{$\tilde{a}_{g\pi(k)}$ {{[answer]}}}\ )\ )$] ) ) where $\pi$ is the random permutation used in $\mathcal{P}_{g}^{\textup{INF}}$, and $\{\tilde{a}_{g\pi(k)}\}_{k=1}^{K}$ are the $K$ numerical values inferred, which form the generated embedding table.
Retrieve embedding from text
Given a generated text instance $\mathcal{G}_{g}\in\mathcal{G}$, we reconstruct the corresponding tabular data by mapping the inferred embedding values $\{\tilde{e}_{g\pi(k)}\}_{k=1}^{K}$ to their respective features. Each textual entry is split into feature-value pairs using "[answer]" as a delimiter. Missing or unrecognized features are assigned the placeholder "NaN". For a specific channel $c$, the output of an inference step $s$ is the reconstructed embedding matrix: $\tilde{E}^{c,s}\in\mathbb{R}^{G\times k_{c}},$ where each row corresponds to a generated instance $\mathcal{G}_{g}$ and each column represents an inferred embedding feature associated with channel $c$. To track all generated embeddings up to step $s$, we define: $\tilde{E}^{c,\leq s}$.
In-generation filtering and stopping criterion
At each inference step $s$, we apply an online filtering procedure that validates generated embeddings without requiring reconstruction into the time-series domain, ensuring efficient real-time evaluation. Specifically, the reconstructed embedding matrices $\big{(}\tilde{E}^{1,s},\dots,\tilde{E}^{C,s}\big{)}$ are filtered based on three criteria: 1) Instances with missing values are discarded, as they prevent accurate reconstruction; 2) Duplicated instances are discarded to maintain diversity in the generated dataset; 3) Significantly diverging instances are discarded (procedure in Appendix A.3). This combined filtering procedure not only enforces diversity and validity among the generated instances but also provides a diagnostic signal: if a substantial fraction of samples is rejected, it may indicate that the fine-tuned LLM requires further training or more representative data.
In generation mode, SDForger employs a dynamic stopping criterion that continues generating batches of $G$ text instances as long as sufficient diversity is preserved among the generated samples (Appendix A.4). However, for consistent comparison with baseline methods across all simulation scenarios, we fix the number of generated instances $\tilde{I}$ across all algorithms. If we denote by $S$ the final inference step, then the output of the generation process is the complete embedding table $\tilde{E}\in\mathbb{R}^{\tilde{I}\times K}$, where, for each channel $c$, $\tilde{E}^{c}=\tilde{E}^{c,\leq S}$.
3.3 Decoding: from tabular embeddings to time series
Given $\tilde{E}$, the time-series representation of generated embeddings can be efficiently recovered due to the reversible nature of the embedding technique used. For the channel $c$, given the generated coefficients $\tilde{e}^{c}_{ij}$ and the corresponding basis system $(b^{c}_{1},\dots,b^{c}_{k_{c}})$, the reconstructed time series are computed as follows Kokoszka and Reimherr (2017): $\tilde{x}^{c}_{i}=\sum_{j=1}^{k_{c}}\tilde{e}^{c}_{ij}b^{c}_{j}$.
This formulation ensures that each generated embedding is decoded back to the original space, resulting in $\tilde{I}$ synthetic instances of a multivariate time series $\tilde{X}_{i}\in\mathbb{R}^{C\times L}$.
4 Evaluation methodology
Evaluation metrics
Evaluating synthetic time-series data requires balancing realism, usability, and efficiency. A strong generative model should replicate key properties of real data while supporting downstream tasks such as forecasting. We adopt a comprehensive evaluation framework comprising two categories: similarity metrics and utility metrics.
  • Similarity metrics, inspired by Ang et al. (2023), assess how closely the generated data matches the real data in terms of distribution, structure, and behavior. They fall into two subtypes: (i) Feature-based metrics which include Marginal Distribution Difference (MDD), Auto-Correlation Difference (ACD), Skewness Difference (SD), and Kurtosis Difference (KD), assess how well synthetic data retains key statistical properties of real data; (ii) Distance-based metrics include Euclidean Distance (ED), Dynamic Time Warping (DTW), and SHAP-RE (SHR), a shapelet-based reconstruction error. They quantify the similarity between synthetic and real data in raw feature space or temporal alignment. Formal definitions are provided in Appendix B.
  • Utility metrics assess the effectiveness of synthetic data in downstream tasks. Specifically, we fine-tune Tiny Time Mixers (TTM) (Ekambaram et al., 2024), a recent foundation model for multivariate time series, under four settings: (1) zero-shot (no fine-tuning) (2) real data only, (3) synthetic data only, and (4) real data augmented with synthetic data. This setup quantifies the impact of synthetic data on model transferability, data efficiency, and robustness.
Evaluation protocols
We consider three distinct evaluation settings to assess the generative capabilities of SDForger across different structural assumptions:
  • Multisample generation aims to produce new instances by combining patterns from multiple existing time series. This setting reflects scenarios such as generating experimental samples, weather profiles, or patient trajectories from heterogeneous observations. It emphasizes diversity and generalization in data-rich contexts.
  • Univariate generation focuses on learning from a single time series to generate plausible alternative versions. This is useful for simulating counterfactual histories, seasonal variations, or stress-test scenarios in domains like finance, weather, and demand forecasting.
  • Multivariate generation evaluates the ability to jointly generate multiple interdependent channels. It reflects real-world settings, such as energy systems, traffic flows, or sensor networks, where channel interactions and cross-correlations are crucial for realism and downstream utility.
In the multisample case, multiple instances are available by design. In contrast, for univariate and multivariate settings, only one instance is provided; therefore, we first apply the period-aware segmentation procedure described in Section 3 to extract multiple windows from each channel.
Parameter settings
We summarize here the hyperparameters for SDForger. We fix the embedding dimension to $\bm{k=3}$ for the multisample and univariate setting. The LLM used for generation is GPT-2  [^1] https://huggingface.co/openai-community/gpt2, fine-tuned with Adam (Diederik, 2014) optimization, a learning rate of $8\times 10^{-5}$, batch size 32, and a maximum of 200 epochs. Early stopping criteria is applied based on the best validation loss computed every 5 steps, patience set to 5, randomly choosing $20\%$ of the data as a validation set.
Baselines
We evaluated SDForger’s performance against several baseline models for synthetic time series generation, covering different approaches. Variational autoencoders: TimeVAE (Desai et al., 2021), which models temporal dependencies with a recurrent VAE architecture, and TimeVQVAE (Lee et al., 2023), which incorporates vector quantization for better capturing discrete temporal patterns; generative adversarial networks: RTSGAN (Pei et al., 2021), which uses recurrent components for adversarial training, and SDEGAN (Kidger et al., 2021), which models time series as solutions to stochastic differential equations; and a diffusion-based model: LS4 (Zhou et al., 2023), which generates sequences via a learned reverse-time diffusion process. Hyperparameters for all baseline competitors follow those reported in their original papers, except for SdeGAN, for which we fix the number of training iterations to $1000$ to balance convergence and computational cost.
Datasets
We evaluated SDForger models using 12 publicly available datasets from various domains, including energy, transport, industry, weather, and finance, with sampling frequencies ranging from 2 minutes to monthly. The datasets were sourced from the Monash Time Series Forecasting Repository and public domains. Detailed information is provided in Appendix C.
5 Results
Following, we discuss results on Similarity-based (Section 5.1), Utility-based metrics (Section 5.2), and a condensed ablation study (Section 5.3). Complete ablations are provided in Appendix D.
5.1 Similarity-based metrics results
The similarity-based results aggregated for the multisample and univariate settings are reported in Table 1, with detailed per-dataset scores provided in Appendix Tables D.7, D.8, D.9, D.10, and D.11.
Overall performance While individual baselines may excel on specific metrics or datasets, SDForger stands out for its consistently strong and balanced performance across all metrics and domains. This pattern is particularly notable for the ICA-based variant, which avoids the common trade-off between statistical and structural fidelity. SDForger’s ability to maintain competitive scores across diverse metrics, as further supported by the average rank scores, underscores its generalization and robustness across heterogeneous temporal datasets. By decoupling representation learning from generation, SDForger is able to capture long-range dependencies while maintaining fidelity to statistical features. This design appears especially effective in producing temporally coherent, domain-consistent samples.
Robustness to evaluation protocols Comparing multisample and univariate settings, we observe that model rankings and relative performances remain largely consistent, suggesting that SDForger is robust to variations in the evaluation protocol. This stability is an important advantage in practice, where test-time conditions may vary.
ICA vs. FPC Between the two embedding strategies, ICA consistently leads, particularly on distance-based metrics. This may reflect the greater interpretability and statistical independence of ICA components, which offer more stable conditioning inputs for generation.
Table 1: Aggregated performance comparison in the multisample and univariate settings. Metrics include raw similarity scores and normalized averages (in $[0-1]$) for each metric group, plus the average rank. Lower values are better. Bold indicates the best performance per column, and underlined indicates the second-best.
Feature-based Distance-based Norm. Avg.
MDD ACD SD KD ED DTW SHR Feat. Dist. Rank
MULTISAMPLE SDF-ICA3 0.244 1.180 0.869 2.384 16.669 12.373 6.870 0.224 0.074 3.143
SDF-FPC3 0.255 2.166 1.323 4.299 17.749 11.921 16.537 0.562 0.100 4.714
TimeVAE 0.227 0.259 0.507 1.697 18.041 11.625 14.021 0.000 0.094 2.143
TimeVQVAE 0.371 5.466 1.327 3.889 13.661 10.167 2.030 0.873 0.000 3.714
RtsGAN 0.279 1.769 0.612 2.300 16.084 11.859 5.631 0.231 0.058 2.857
SdeGAN 0.240 2.098 1.404 4.091 37.174 33.391 51.678 0.540 0.693 5.286
LS4 0.276 6.150 1.243 4.852 44.389 31.806 160.403 0.789 0.977 6.143
UNIVARIATE SDF-ICA3 0.306 1.396 0.671 1.382 18.802 12.435 4.856 0.149 0.070 2.429
SDF-FPC3 0.308 1.480 0.801 1.690 19.340 12.809 5.452 0.354 0.084 4.000
TimeVAE 0.288 2.013 0.611 1.245 20.778 12.126 18.534 0.066 0.158 2.714
TimeVQVAE 0.433 4.330 0.740 2.052 15.438 11.250 2.217 0.707 0.000 3.571
RtsGAN 0.363 2.389 0.776 1.325 18.951 12.926 5.464 0.384 0.081 4.000
SdeGAN 0.267 3.659 0.813 1.542 42.017 38.541 65.557 0.390 0.979 5.143
LS4 0.298 6.041 0.855 2.457 40.362 24.262 69.751 0.797 0.805 6.143
5.2 Utility-based metrics results
Table 2: Utility evaluation via fine-tuned forecasting models. TTM forecasting performance on downstream tasks using different training sources: zero-shot, original data, generated data, and a combination of original and generated data. Results are reported for 3 multivariate datasets: bikesharing (target: count, control: temperature, humidity), etth1 (target: HUFL, control: MUFL, OT), and traffic (target: junction1, control: junction2, junction3). Metrics include RMSE, MASE, WQL, and average rank (lower is better). Bold highlights the best result within each row group; underlined the second best; bold+underlined the overall best.
bikesharing etth1 traffic
RMSE MASE WQL RMSE MASE WQL RMSE MASE WQL Avg. Rank
0-shot 0.728 2.150 0.287 0.678 2.132 0.255 0.708 1.555 0.255 1.78
Original Data (OD) 0.495 0.822 0.178 0.658 1.820 0.232 0.702 1.995 0.283 1.22
GENERATED SDF-ICA 0.514 0.899 0.194 0.626 1.820 0.224 0.655 1.849 0.262 2.00
SDF-FPC 0.527 0.926 0.200 0.650 1.887 0.232 0.662 1.837 0.262 3.22
TimeVAE 0.566 0.983 0.211 0.690 2.268 0.269 0.738 2.078 0.296 5.33
TimeVQVAE 0.520 0.867 0.188 0.626 1.874 0.227 0.702 1.995 0.283 2.67
RtsGAN 0.710 1.261 0.275 0.770 2.271 0.291 0.597 1.574 0.225 4.67
SdeGAN 0.572 0.995 0.214 0.688 2.262 0.263 0.629 1.715 0.243 4.00
LS4 0.839 1.468 0.318 0.642 1.977 0.236 0.917 2.595 0.369 5.89
ORIGINAL + GEN SDF-ICA + OD 0.487 0.801 0.173 0.642 1.746 0.226 0.750 2.110 0.301 3.22
SDF-FPC + OD 0.493 0.829 0.179 0.666 1.754 0.231 0.743 2.087 0.297 4.78
TimeVAE + OD 0.492 0.814 0.176 0.654 1.752 0.228 0.721 2.039 0.290 2.89
TimeVQVAE + OD 0.495 0.804 0.174 0.678 1.887 0.242 0.724 2.043 0.291 4.56
RtsGAN + OD 0.498 0.819 0.177 0.637 1.872 0.231 0.607 1.647 0.234 3.44
SdeGAN + OD 0.495 0.837 0.181 0.620 1.843 0.224 0.605 1.716 0.242 3.33
LS4 + OD 0.497 0.822 0.178 0.660 1.819 0.233 0.745 2.111 0.300 5.56
Table 2 presents the utility evaluation, where we assess the practical value of synthetic data by fine-tuning TTM on different multivariate training sources. For LS4, we modified the architecture to support multivariate generation. We also fine-tune the embedding dimension $k$ in SDForger based on dataset complexity: $k{=}3$ for bikesharing, $k{=}7$ for etth1, and $k{=}5$ for traffic.
SDForger demonstrates strong performance across datasets, as evidenced by its top average rank, with notable results on bikesharing and etth1. In bikesharing, synthetic data from SDForger alone yields competitive scores, and combining it with real data leads to the best overall performance across metrics. On etth1, SDForger-generated data surpasses original data in RMSE and WQL, suggesting it captures critical temporal and statistical structure. The hybrid setting (original + generated) maintains this advantage and further improves MASE. Performance on traffic is more nuanced. Here, fine-tuning on real data is less effective, and GAN-based methods outperform others. Nevertheless, SDForger remains competitive, especially when using synthetic data alone. This suggests that the test distribution may deviate significantly from the training set, making traditional fine-tuning less useful. Indeed, high-quality synthetic data can act as a valuable supplement or even an alternative.
In no scenario does synthetic data degrade downstream performance, underscoring the reliability and utility of SDForger-generated samples across varied forecasting contexts.
5.3 Ablation
Effect of embedding dimension $\bm{k}$ Appendix Table D.4 presents an ablation study on the number of components $k$ used in SDForger’s embedding space. A compact embedding with $k{=}3$ offers strong performance across both multisample and univariate settings, indicating that a small number of components is often sufficient to capture core temporal and structural patterns. However, the optimal value of $k$ may vary across datasets, with more intricate dynamics potentially requiring higher-dimensional representations. In practice, users may also opt to select $k$ based on a desired percentage of explained variance, adapting the representation to specific application needs.
Domain-level insights Appendix Tables D.1 and D.2 summarize the average normalized similarity scores per dataset. SDForger models achieve strong performance across a wide range of domains. Structured datasets such as Energy, Appliances, and Weather exhibit particularly high scores. In contrast, domains such as Tourism, Traffic, and Finance present greater challenges, likely due to their increased irregularity and noise. Nonetheless, SDForger maintains competitive results even in more complex settings, underscoring the flexibility of the proposed architecture.
Generation efficiency Appendix Table D.3 reports the average time to generate univariate sequences for three targets from the Bikesharing dataset, across two window lengths. SDForger is substantially faster than all competitors, often by one to two orders of magnitude. TimeVAE is the closest competitor but remains over 4$\times$ slower. Notably, unlike GAN-based competitors, SDForger’s generation time is independent of sequence length and scales with the number of embedding components ($k$). A minor exception occurs at $k=3$, where the reduced latent expressivity increases LLM fine-tuning time. Overall, SDForger achieves state-of-the-art efficiency without compromising quality.
LLM comparison GPT-2 (124M) achieves performance on par with, and sometimes better than larger and more recent models such as granite-3.0 (2B) and phi-3.5 (3.8B) (Appendix Table D.6). This shows that SDForger’s pipeline is effective even leveraging lightweight models. While runtime cost grows with model size, SDForger remains efficient compared to baselines (Appendix Table D.5).
6 Shaping time series with language
SDForger is designed to naturally incorporate textual information, making it well-suited for state-of-the-art time-series generation that embraces additional multi-modal inputs. To explore this, we conduct an experiment using the bikesharing dataset. Coming from the intuition that these variables stem from a common physical process and may share latent components, we embed their three channels (temperature, count, and humidity) into a shared ICA basis. We incorporate the channel information in the textual encoder: ‘‘Condition: data is temp [sep] Input: value_1 is [blank]...[sep] Target: $e_{i1}$ [answer]...’’
This conditioning strategy enables SDForger to generate channel-specific sequences with high fidelity. For instance, using a longitudinal $k$-nearest neighbor classifier (Ramos-Carreño et al., 2024) trained on real data, we achieve an accuracy of $0.81$ in identifying the generated curves (see Figure 2). These results highlight SDForger’s strong generative capacity and its ability to integrate and respond to textual cues, positioning it as a flexible and powerful baseline for multimodal time-series synthesis.
Figure 2: Text-Conditioned Generation with SDForger. Visualization of 10 original (grey) and synthetic samples per channel from the bikesharing data. Synthetic data is generated using conditional prompts: “Condition: data is cnt (blue)”, “Condition: data is hum (pink)”, and “Condition: data is temp (orange)”.
7 Conclusions
We introduced SDForger, a flexible and efficient framework for generating synthetic multivariate time series using large language models. By combining compact functional embeddings with textual conditioning, SDForger enables high-quality generation even in data-scarce settings. Extensive evaluations across multiple datasets and tasks demonstrate that SDForger consistently achieves strong similarity scores and enhances downstream forecasting performance—often matching or surpassing results obtained from real data and outperforming state-of-the-art baselines.
Ablation studies confirm the robustness of the framework across embedding strategies, dimensionality choices, and LLM architectures. SDForger is also highly efficient, with significantly lower generation times compared to its competitors. Moreover, by leveraging LLMs, SDForger enables seamless integration with textual prompts, paving the way for multimodal time-series generation, where natural language can guide not only content but also structure, semantics, or temporal context.
One limitation of our approach is the trade-off between embedding dimensionality and training set size: using too many components relative to the number of instances can lead to unstable fine-tuning and reduced generation quality. Additionally, our current implementation supports only autoregressive LLMs; future work could extend the framework to encoder-only models and different generation paradigms such as masked token prediction. Further improvements may also come from adaptive embedding strategies and adding contextual information (Section 6) for more reliable generation.
Appendix
Appendix A Implementation details
A.1 Data preprocessing: segmentation
In many scenarios, each channel consists of a single historical time series, i.e., $I_{0}=1$. However, to estimate embeddings that effectively capture the temporal distribution, multiple instances per channel are necessary. Therefore, when $I_{0}=1$, we segment each channel into multiple overlapping windows. Specifically, for each channel $c$, we construct $I_{c}$ windows of fixed length $L_{c}$, where $L_{c}<L_{0}$. Without loss of generality, for simplicity in notation, we assume $L_{c}$ and $I_{c}$ are identical across all channels and denote them as $L$ and $I$, respectively.
To ensure robust learning of the embedding distribution, $I$ must be sufficiently large. Our experiments indicate that even $I=15$ (i.e., 15 instances) suffices for this purpose. Once $I$ and $L$ are fixed, we segment the time series while minimizing the overlap between consecutive windows. The overlap step is determined by the dominant periodicity $P$ of the channel, ensuring that window transitions align with intrinsic temporal cycles.
The set of extracted windows $\mathcal{W}$ is formally defined as:
$\mathcal{W}=\{X[t:t+L]\mid t=0,s,2s,\dots,L_{0}-L\}$ (1)
where the step size $s$ is computed as: $s=\max(1,\lfloor\frac{L_{0}-L}{I-1}\rfloor)$ and then adjusted to be the nearest multiple of $P$ to maintain consistency in periodic structure.
To determine the dominant periodicity $P$, we employ the Autocorrelation Function (ACF), which quantifies the similarity between the time series and its lagged versions at different time shifts. This method is robust to noise and remains effective even when periodicity is not strictly stationary.
The estimation of $P$ follows these steps:
  1. 1.
    Compute the ACF and identify significant peaks, excluding lag 00.
  2. 2.
    Rank the detected peaks by their autocorrelation values.
  3. 3.
    Select the highest-ranked period $P$ such that $P<L/2$.
Figure A.1: Periodicity-aware segmentation.
By leveraging this periodicity-aware segmentation strategy, we ensure that the extracted windows align with the natural cycles of the time series. Moreover, this approach minimizes window overlap, maximizing their independence and diversity and facilitating more effective embedding computations for downstream generative tasks.
This pre-processing transforms the data from a single sequence $X_{1}$ into a set $X=\{X_{i}\}_{i=1}^{I}$, where each $X_{i}\in\mathbb{R}^{C\times L}$, preparing the data for the generation task.
A.2 Choice of the number of components
The choice of $k_{c}$ determines how well the basis representation approximates the original time-series channel $c$. Our frameword allows the user to either select the smallest $k_{c}$ that explains a predefined percentage of the total variance of the original time series or manually specify $k_{c}$.
There exists an inherent trade-off in selecting $k_{c}$. A higher $k_{c}$ captures more of the total variability but can hinder the LLM’s ability to model the underlying distribution during the generation phase. Moreover, if $k_{c}$ is too large, the generated samples may become overly similar to the original data, limiting diversity and the introduction of novel patterns. Conversely, choosing $k_{c}$ too small risks omitting essential structures and temporal characteristics, degrading the reconstruction quality.
It is important to note that the nature of the components differs between FPC and FastICA. FPC forms a parsimonious basis system, where a few components typically suffice to capture most of the variability, with components ordered by the amount of variance they explain—early components being systematically more informative. In contrast, FastICA components are unordered: each component contributes independently, without a hierarchical importance structure. As a result, FastICA generally requires more components to achieve a similar reconstruction quality compared to FPC. However, this property also makes FastICA embeddings more robust during generation, as information is distributed more evenly across components, reducing the risk that a few badly generated components disproportionately affect the synthesized curves.
In practice, we recommend keeping the total number of components $K$ reasonably small, particularly when the number of training instances is limited. Empirically, with a training set of 30 instances, setting $K>25$ often results in unstable fine-tuning and an increased rate of discarded samples due to low-quality generation. This limitation stems from the LLM’s reduced ability to effectively model high-dimensional embeddings under data-scarce conditions.
A.3 In-generation filtering
To ensure the quality of synthetic data, we discard generated instances whose embedding coefficients significantly deviate from the distribution of the original data. Specifically, we compute the squared $\ell_{2}$-norm of each embedding vector and compare it to the norms of the original embeddings. This criterion efficiently filters out extreme outliers in the latent space, without requiring reconstruction into the time-series domain.
Formally, for each channel $c$, let $\hat{E}^{c,s}=E^{c}\cup\tilde{E}^{c,\leq s-1}$ be the matrix containing both original embeddings and all previously accepted generated embeddings up to inference step $s-1$, with $\hat{E}^{c,0}=E^{c}$. Denote by $N^{c,\text{old}}$ and $N^{c,\text{new}}$ the sets of squared Euclidean norms of the rows of $\hat{E}^{c,s-1}$ and the newly generated matrix $\tilde{E}^{c,s}$, respectively. For each newly generated row $i$, we compute its norm and accept it only if:
$q_{1}-3\cdot\text{IQR}\leq N^{c,\text{new}}_{i}\leq q_{3}+3\cdot\text{IQR},$
where $q_{1}$ and $q_{3}$ are the first and third quartiles of $N^{c,\text{old}}$, and $\text{IQR}=q_{3}-q_{1}$. An instance is retained only if this condition is satisfied across all channels $c$.
This norm-based strategy is particularly well-motivated when using FPCs, due to the orthonormality of the basis. Let $\mathcal{X}^{c}_{i}$ be a time series in channel $c$ and $b^{c}_{j}$ the corresponding FPC basis. Then the $\mathbb{L}^{2}$-norm of $\mathcal{X}^{c}_{i}$ can be approximated by the Euclidean norm of its FPC coefficients:
$\int_{\mathcal{T}}(\mathcal{X}^{c}_{i})^{2}\,\mathrm{d}t=\lVert\mathcal{X}^{c}_{i}\rVert_{\mathbb{L}^{2}}^{2}=\sum_{j=1}^{\infty}\langle\mathcal{X}^{c}_{i},b^{c}_{j}\rangle_{\mathbb{L}^{2}}^{2}\approx\sum_{j=1}^{k_{c}}(e^{c}_{ij})^{2},$
where $e^{c}_{ij}$ are the FPC embedding coefficients. This justifies the norm-based filtering as a direct proxy for detecting time-series samples with unusually high or low energy.
While filtering based on coefficient norms does not guarantee full statistical fidelity to the original time-series distribution, it serves as an effective mechanism to remove extreme outliers without reconstruction. Combined with additional checks for missing values and duplicates, this step helps preserve both the diversity and relevance of the generated data. A high rejection rate may indicate insufficient LLM fine-tuning or poor generalization, suggesting the need for more representative training data or additional training steps.
A.4 Stopping criterion
The stopping criterion monitors the diversity of the generated norms across all channels to determine when the generation process should stop. When using FPC, these norms correspond to the $\mathbb{L}^{2}$-norms of the generated curves. When using FastICA, they correspond to the norms of the embedding coefficient vectors; although not directly related to the curve norms, they still provide a useful proxy for identifying over-sampling and loss of variability.
At inference step $s$, for each channel $c$, let $u^{c}$ denote the number of unique values in $N^{c,\text{old}}$ (the set of accepted norms up to step $s$), rounded to the fourth decimal place. Let $\tilde{I}$ be the total number of valid instances generated so far. We define the diversity score for channel $c$ as:
$D^{c}={u^{c}}/{\tilde{I}}.$
The diversity score provides a quantitative measure of how much variability remains in the generated norms. We track $D^{c}$ at each inference step, and the stopping condition is triggered when:
$\max_{c}D^{c}<\lambda_{\text{stop}}\quad\text{or}\quad\tilde{I}>\tilde{I}_{\max}.$
In other words, generation stops either when the maximum diversity score across channels falls below a predefined threshold $\lambda_{\text{stop}}$, indicating reduced novelty, or when the total number of generated instances exceeds a maximum cap $\tilde{I}_{\max}$.
Monitoring the diversity score enables us to assess whether the model continues to introduce new variability in the generated data, serving as an online signal for generation quality.
If we denote by $S$ the final inference step, then the output of the generation process is the complete embedding table $\tilde{E}\in\mathbb{R}^{\tilde{I}\times K}$, where, for each channel $c$, $\tilde{E}^{c}=\tilde{E}^{c,\leq S}$.
Appendix B Evaluation metrics
All the metrics presented below are adopted from Ang et al. (2023), except for Shapelet-based Reconstruction, which follows the definition in Zheng et al. (2016).
B.1 Feature-based evaluation
Marginal Distribution Difference.
MDD computes an empirical histogram for each dimension and time step in the generated series, using the bin centers and widths from the original series. It then calculates the average absolute difference between this histogram and that of the original series across bins, assessing how closely the distributions of the original and generated series align.
AutoCorrelation Difference.
ACD computes the autocorrelation of both the original and generated time series, then determines their difference. By contrasting the autocorrelations, we could evaluate how well dependencies are maintained in the generated time series.
Skewness Difference.
SD is vital for the marginal distribution of a time series, quantifying its distribution asymmetry. Given the mean (standard deviation) of the train time series $T_{s}^{tr}$ as $\mu_{s}^{tr}$ ($\sigma_{s}^{tr}$) and the generated time series $T_{s}^{gen}$ as $\mu_{s}^{gen}$ ($\sigma_{s}^{gen}$), we evaluate the fidelity of $T_{s}^{gen}$ by computing the skewness difference between them as:
$SD=\left|\frac{\mathbb{E}[(T_{s}^{gen}-\mu_{s}^{gen})^{3}]}{\sigma_{s}^{gen3}}-\frac{\mathbb{E}[(T_{s}^{tr}-\mu_{s}^{tr})^{3}]}{\sigma_{s}^{tr3}}\right|.$
Kurtosis Difference.
Like skewness, KD assesses the tail behavior of a distribution, revealing extreme deviations from the mean. Using the previous notations, the kurtosis difference between $T_{s}^{tr}$ and $T_{s}^{gen}$ is calculated as:
$KD=\left|\frac{\mathbb{E}[(T_{s}^{gen}-\mu_{s}^{gen})^{4}]}{\sigma_{s}^{gen4}}-\frac{\mathbb{E}[(T_{s}^{tr}-\mu_{s}^{tr})^{4}]}{\sigma_{s}^{tr4}}\right|.$
B.2 Distance-based evaluation
Euclidean Distance.
For each original series $s^{tr}=(x_{1},...,x_{l})$ and its generated $s^{gen}=(y_{1},...,y_{l})$, $ED=\sqrt{\sum_{1=1}^{l}(x_{i}-y_{i})^{2}}$. We take the mean of ED for all series and all samples. Given that the input time series has been preprocessed to fit within the range of $[0,1]$, ED deterministically assesses the similarity between $s^{gen}$ and $s^{tr}$. It provides a value-wise comparison between the time series.
Dynamic Time Warping.
Given that ED overlooks alignment, we include DTW to capture the optimal alignment between series regardless of their pace or timing. The alignment facilitated by DTW offers insights into the predictive quality of the generated series.
Shapelet-based Reconstructions.
Shapelet based RE is calculated by generated time series using shapelets extracted using shift invariant dictionary learning (SIDL) algorithm (Zheng et al., 2016). Shapelets represent local discriminative patterns present in the time-series data. We learn shift invariant patterns/shapelets on the original time-series dataset and then use the learnt dictionary to reconstruct unseen generated time series. The reconstruction error is calculated between the generated time series and their reconstruction using SIDL.
Appendix C Datasets & protocol settings
Table C.1: Overview of the benchmark datasets. For each dataset, we report its application domain, sampling frequency, number of time series, length statistics, and the type of evaluation (univariate, multivariate, multi-sample) it supports.
Dataset Domain Freq. Number Evaluation type
MS UV MV
Australian Electricity energy 30m 5 N Y Y
Appliances energy 10m 1 N Y N
Bikesharing general 1H 3 N Y Y
Carbon Capture Plant nature 2m 4 N Y N
ETTH1 energy 1H 3 N Y Y
ECL energy 1H 320 Y N N
Exchange Rate finance 1D 8 N Y N
NN5 finance 1D 111 Y N N
Tourism general 1M 365 Y N N
Traffic transport 1H 3 N Y Y
Traffic Monash transport 1H 861 Y N N
Solar - Weather nature 1H 653 Y N N
Rain - Weather nature 1H 386 Y N N
Temperature - Weather nature 1H 362 Y N N
C.1 Dataset overview
Energy
Mobility and Transport
Nature
  • Carbon Capture Plant (Jablonka et al., 2023) records the emission profiles of “2-amino-2-methyl-1-propanol” (AMP) and “piperazine” (Pz) collected at every 2 minutes interval.
  • Weather (Godahewa et al., 2021) contains daily time series of four weather variables (rain, mintemp, maxtemp and solar radiation) measured at weather stations in Australia.
Finance
  • Exchange Rate (Godahewa et al., 2021) contains daily exchange rates for currencies of eight countries (Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore) between 1990 and 2016.
  • NN5 (Daily, Weekly) (Godahewa et al., 2021) contains cash withdrawal data from ATMs.
C.2 Protocols
Multisample setting
For MS data preparation, we sampled $I=30$ instances, each of length $L=250$, resulting in a total training sequence of 15,000 timestamps. Standard scaling per timestamp is applied. For evaluation, we generate 100 synthetic instances.
Univariate setting
For UV data preparation, we used a training sequence of length $L_{0}=2,000$, segmented into $I=30$ instances of length $L=250$ using our periodicity-aware segmentation strategy (cf Appendix A.1). Same standard scaling was applied. For evaluation, we generate 100 synthetic instances.
Multivariate setting
For fine-tuning TTM in the MV setting, we used consecutive sequences of length $L_{0}=5000$, $2500$, and $2500$ for training, validation, and testing, respectively. Each set was segmented into $I=30$, $15$, and $15$ instances of length $L=1120$ using our period-aware segmentation strategy (Appendix A.1). Periodicity was estimated from the training set. Standard scaling per timestamp, computed on the training set, was applied consistently across all splits. For evaluation, we generated 30 instances, matching the size of the training set.
Appendix D Additional results
Table D.1: Per-dataset similarity results in the multisample setting. Average normalized similarity scores (feature-based and distance-based) for each dataset and model.
Feature-Based Distance-Based
ecl nn5 tourism traffic weather ecl nn5 tourism traffic weather
SDF-ICA3 0.402 0.195 0.268 0.330 0.204 0.018 0.129 0.032 0.086 0.137
SDF-FPC3 0.576 0.323 0.594 0.293 0.296 0.054 0.183 0.073 0.099 0.136
TimeVAE 0.164 0.109 0.211 0.143 0.174 0.063 0.139 0.074 0.143 0.125
TimeVQVAE 0.754 0.380 0.818 0.437 0.498 0.003 0.068 0.009 0.053 0.069
RtsGAN 0.049 0.347 0.174 0.344 0.315 0.075 0.074 0.110 0.059 0.102
SdeGAN 0.499 0.219 0.539 0.391 0.308 0.299 0.652 0.290 0.496 0.616
LS4 0.860 0.319 0.849 0.470 0.400 0.720 0.681 0.639 0.926 0.707
Table D.2: Per-dataset similarity results in the univariate setting. Average normalized similarity scores (feature-based and distance-based) for each dataset and model.
Feature-Based Distance-Based
appl. austr. bike carbon etth1 exch. traffic appl. austr. bike carbon etth1 exch. traffic
SDF-ICA3 0.486 0.122 0.162 0.197 0.128 0.155 0.269 0.088 0.073 0.077 0.063 0.080 0.115 0.059
SDF-FPC3 0.509 0.122 0.169 0.345 0.107 0.287 0.213 0.085 0.101 0.077 0.078 0.076 0.143 0.064
TimeVAE 0.370 0.123 0.184 0.176 0.234 0.113 0.236 0.119 0.079 0.155 0.102 0.140 0.091 0.178
TimeVQVAE 0.531 0.439 0.371 0.572 0.354 0.578 0.312 0.045 0.031 0.037 0.005 0.031 0.031 0.035
RtsGAN 0.567 0.300 0.273 0.226 0.244 0.217 0.312 0.056 0.098 0.084 0.085 0.072 0.157 0.053
SdeGAN 0.634 0.128 0.174 0.409 0.173 0.115 0.249 0.456 0.623 0.831 0.703 0.833 0.469 0.765
LS4 0.609 0.294 0.265 0.695 0.289 0.416 0.300 0.720 0.471 0.475 0.554 0.364 0.498 0.621
Table D.3: Average generation time: baselines Average time (in seconds) required to generate synthetic univariate time series for the bikesharing dataset across three targets: count, temperature, and humidity. We report results for two input sequence lengths: 250 and 500. All models were evaluated under the same computational constraints (-mem 20G -cores 1+1 -gpu v100) using a single NVIDIA V100 GPU.
Length SDF-ICA3 SDF-ICA5 SDF-ICA7 SDF-FPC3 SDF-FPC5 SDF-FPC7 TimeVAE TimeVQVAE RtsGAN SdeGAN LS4
250 41.9 26.8 28.8 22.0 25.4 33.1 138.1 4574.2 2055.7 3498.9 2804.4
500 38.3 22.8 26.0 17.9 22.8 26.6 112.6 4401.4 3536.3 7316.7 2378.9
Table D.4: Ablation study: embedding dimension. Aggregated similarity-based performance across all datasets in the multisample and univariate setting.
Feature-based Distance-based Norm. Avg.
MDD ACD SD KD ED DTW SHR Feat. Dist.
MULTISAMPLE SDF-FPC3 0.255 2.166 1.323 4.299 17.749 11.921 16.537 0.616 0.609
SDF-FPC5 0.262 3.191 1.336 3.668 17.475 11.727 22.893 0.714 0.535
SDF-FPC7 0.264 3.534 1.500 3.560 17.710 11.652 28.068 0.787 0.655
SDF-ICA3 0.244 1.180 0.869 2.384 16.669 12.373 6.870 0.050 0.333
SDF-ICA5 0.261 0.782 1.378 2.649 16.743 12.238 7.731 0.371 0.307
SDF-ICA7 0.265 0.589 1.964 2.963 16.900 12.031 14.195 0.576 0.362
UNIVARIATE SDF-FPC3 0.308 1.480 0.801 1.690 19.340 12.809 5.452 0.736 0.469
SDF-FPC5 0.306 1.887 0.773 1.581 20.534 12.513 8.920 0.536 0.753
SDF-FPC7 0.309 2.399 0.774 1.954 20.470 12.079 10.982 0.947 0.654
SDF-ICA3 0.306 1.396 0.671 1.382 18.802 12.435 4.856 0.169 0.163
SDF-ICA5 0.306 0.867 0.770 1.333 19.043 12.261 6.555 0.279 0.222
SDF-ICA7 0.306 0.597 0.736 1.458 19.989 12.381 8.102 0.175 0.543
Table D.5: Average Generation Time Across LLM Backbones. Average time (in seconds) required to generate synthetic univariate time series for the bikesharing dataset across three targets: count, temperature, and humidity. We report results for two input sequence lengths (250 and 500) and compare three LLM backbones: GPT-2, granite-3.0-2b-base, and Phi-3.5-mini-instruct. All models were evaluated under the same computational constraints (-mem 100G -cores 1+1 -gpu a100) using a single NVIDIA A100 GPU. For fine-tuning, we use a batch size of 16 for granite and 8 for phi.
Length ICA3 + gpt2 ICA3 + granite ICA3 + phi ICA5 + gpt2 ICA5 + granite ICA5 + phi ICA7 + gpt2 ICA7 + granite ICA7 + phi
250 22.7 112.8 132.6 18.3 118.5 113.9 19.0 119.9 126.1
500 16.2 93.6 98.5 17.3 99.9 103.7 18.9 125.0 110.7
Table D.6: Ablation study: LLM backbone. Aggregated similarity-based performance across all datasets for different LLMs used in SDF models. We compare GPT-2 with two larger and more recent alternatives: granite-3.0-2b-base [^9] https://huggingface.co/ibm-granite/granite-3.0-2b-base (2B parameters) and Phi-3.5-mini-instruct [^10] https://huggingface.co/microsoft/Phi-3.5-mini-instruct (3.8B parameters). For fine-tuning, we use a batch size of 16 for granite and 8 for phi.
Feature-based Distance-based Norm. Avg.
MDD ACD SD KD ED DTW SHR Feat. Dist.
MULTISAMPLE SDF-FPC3 + GPT-2 0.255 2.166 1.323 4.299 17.749 11.921 16.537 0.964 0.747
SDF-FPC3 + Granite 0.251 1.817 1.227 4.132 16.429 11.659 11.565 0.757 0.245
SDF-FPC3 + Phi-3 0.257 1.215 1.154 3.723 16.734 11.872 12.367 0.643 0.397
SDF-ICA3 + GPT-2 0.244 1.180 0.869 2.384 16.669 12.373 6.870 0.101 0.361
SDF-ICA3 + Granite 0.241 1.069 0.961 2.524 16.953 12.744 6.160 0.101 0.509
SDF-ICA3 + Phi-3 0.247 0.907 1.102 3.570 16.069 11.847 6.499 0.382 0.069
UNIVARIATE SDF-FPC3 + GPT-2 0.308 1.480 0.801 1.690 19.340 12.809 5.452 0.947 0.804
SDF-FPC3 + Granite 0.305 1.268 0.673 1.185 19.026 12.556 5.457 0.207 0.505
SDF-FPC3 + Phi-3 0.310 1.368 0.671 1.305 19.561 12.767 5.196 0.574 0.777
SDF-ICA3 + GPT-2 0.306 1.396 0.671 1.382 18.802 12.435 4.856 0.496 0.161
SDF-ICA3 + Granite 0.304 1.370 0.679 1.123 18.616 12.365 4.712 0.253 0.000
SDF-ICA3 + Phi-3 0.307 1.398 0.541 1.199 19.081 12.471 5.856 0.337 0.577
Table D.7: Multisample evaluation: similarity metrics reported per dataset.
SDForger Models VAE Models GAN Models Others
ICA3
FPC3
TimeVAE
TimeVQVAE
RTSGAN
SDEGAN
LS4
ECL
MDD ACD SD KD ED DTW SHAP-RE 0.154 0.146 2.47 7.342 10.141 9.715 0.424 0.219 4.492 2.762 6.07 12.543 10.985 4.027 0.156 0.082 0.746 3.294 12.887 12.025 1.922 0.292 7.862 2.803 6.895 9.978 7.97 2.21 0.145 0.051 0.114 1.001 13.461 12.486 4.123 0.193 0.174 2.91 8.673 25.951 25.431 7.048 0.296 8.365 2.983 10.074 43.715 40.142 105.287
NN5
MDD ACD SD KD ED DTW SHAP-RE 0.248 1.489 0.307 1.43 19.308 12.837 9.482 0.248 3.235 0.428 4.249 20.576 12.617 40.975 0.243 0.221 0.126 0.151 20.514 11.223 20.993 0.371 4.964 0.259 1.159 15.019 10.914 2.072 0.383 4.646 0.092 0.348 16.201 9.684 8.254 0.246 2.73 0.422 0.512 43.433 36.712 83.918 0.262 5.677 0.287 0.96 38.822 24.415 207.419
Tourism
MDD ACD SD KD ED DTW SHAP-RE 0.189 0.22 1.321 4.477 11.216 10.291 0.547 0.24 3.29 2.613 8.297 14.039 11.837 4.28 0.172 0.206 0.854 4.251 13.895 12.409 1.785 0.339 7.807 2.896 7.877 10.516 8.192 2.135 0.121 0.272 0.681 4.852 15.405 14.741 2.833 0.208 0.215 2.988 9.616 25.399 25.098 6.409 0.282 8.228 2.806 10.907 39.897 35.521 100.231
Traffic
MDD ACD SD KD ED DTW SHAP-RE 0.251 1.443 1.433 3.177 16.532 10.917 6.568 0.234 1.368 1.507 1.937 18.429 10.5 9.188 0.234 0.097 0.353 1.263 20.748 12.343 14.948 0.359 3.767 1.377 1.503 14.169 10.039 1.995 0.314 0.886 1.384 2.642 15.552 9.164 5.101 0.222 3.749 1.598 3.113 35.522 31.69 51.49 0.242 5.229 1.403 4.727 55.908 37.028 205.07
Weather
MDD ACD SD KD ED DTW SHAP-RE 0.292 1.097 0.131 0.419 25.51 20.532 6.638 0.293 2.221 0.017 2.561 21.661 15.354 24.439 0.282 0.533 0.435 0.591 19.665 13.238 9.079 0.447 6.717 0.005 2.188 15.113 11.219 2.098 0.303 1.15 0.37 0.584 19.77 15.884 3.931 0.286 0.705 0.237 0.527 43.035 41.331 34.196 0.296 7.597 0.142 0.976 46.189 34.55 152.274
Weather
MDD ACD SD KD ED DTW SHAP-RE 0.271 1.705 0.185 1.065 17.199 13.148 6.217 0.27 1.053 0.216 1.789 18.533 11.255 14.613 0.264 0.278 0.304 0.757 19.575 10.48 15.605 0.41 5.937 0.05 1.936 15.177 10.675 1.956 0.365 1.508 0.398 0.985 15.575 10.837 2.587 0.266 1.67 0.249 0.213 44.094 41.651 55.846 0.275 6.737 0.145 1.174 46.266 32.415 208.289
Weather
MDD ACD SD KD ED DTW SHAP-RE 0.233 1.642 0.599 1.146 14.461 10.935 10.955 0.222 0.321 2.474 6.636 16.009 11.493 9.555 0.175 0.309 1.065 3.238 16.979 11.35 25.464 0.291 2.491 2.683 7.94 13.631 10.718 1.883 0.259 2.735 1.641 7.447 15.033 10.803 9.595 0.199 5.549 2.729 9.639 31.412 28.891 57.118 0.27 4.696 2.135 9.724 50.966 33.017 224.024
Weather
MDD ACD SD KD ED DTW SHAP-RE 0.314 1.695 0.508 0.013 18.989 10.61 14.132 0.31 1.351 0.564 2.854 20.201 11.322 25.218 0.29 0.349 0.172 0.029 20.061 9.932 22.375 0.459 4.18 0.539 1.618 15.69 11.607 1.892 0.342 2.902 0.216 0.544 17.677 11.275 8.623 0.297 1.996 0.101 0.433 48.549 36.321 117.395 0.284 2.667 0.044 0.27 33.346 17.363 80.632
Table D.8: Univariate evaluation: similarity metrics reported for Energy datasets.
SDForger Models VAE Models GAN Models Others
ICA3
FPC3
TimeVAE
TimeVQVAE
RTSGAN
SDEGAN
LS4
Appliances
MDD ACD SD KD ED DTW SHAP-RE 0.318 1.825 2.068 3.249 19.493 11.597 9.588 0.315 1.714 2.093 3.816 19.13 11.73 9.072 0.303 3.137 1.008 2.891 20.459 10.03 27.262 0.405 1.714 2.07 2.586 15.916 12.322 1.952 0.414 2.404 2.152 2.559 17.162 11.301 6.367 0.241 7.208 2.269 4.196 29.562 27.191 71.477 0.251 4.301 2.028 5.874 47.81 16.564 162.749
Australian Elec
MDD ACD SD KD ED DTW SHAP-RE 0.315 1.069 0.165 0.161 19.017 10.652 3.411 0.312 0.392 0.103 0.452 22.663 13.655 2.719 0.292 1.431 0.65 1.296 19.656 10.899 7.925 0.457 5.641 0.071 2.12 15.657 10.583 2.441 0.376 1.926 0.3 1.829 22.511 14.942 4.007 0.281 2.976 0.091 0.325 42.67 40.842 23.325 0.323 7.129 0.423 0.765 38.681 26.104 34.777
Australian Elec
MDD ACD SD KD ED DTW SHAP-RE 0.287 1.46 0.274 1.8 18.368 12.248 2.211 0.281 0.457 0.307 1.677 19.769 12.842 2.91 0.277 0.265 0.11 0.503 19.51 10.312 6.061 0.436 5.579 0.137 1.99 15.717 10.648 2.587 0.421 4.796 0.589 0.264 20.145 13.322 3.291 0.273 2.518 0.031 0.273 44.034 41.201 26.765 0.293 6.908 0.032 0.899 37.357 25.029 30.768
Australian Elec
MDD ACD SD KD ED DTW SHAP-RE 0.313 1.196 0.42 0.179 20.624 13.17 2.683 0.313 0.569 0.718 0.69 21.174 13.104 3.04 0.292 0.237 0.322 0.539 19.537 10.626 4.984 0.486 5.675 0.598 2.111 15.889 10.748 2.38 0.334 2.541 0.605 2.386 19.896 14.559 1.572 0.297 2.559 0.387 0.389 45.934 44.425 23.29 0.315 7.209 0.229 0.804 37.731 26.567 28.057
Australian Elec
MDD ACD SD KD ED DTW SHAP-RE 0.283 1.549 0.399 0.057 17.406 11.621 1.888 0.289 1.35 0.389 0.465 19.377 14.456 1.929 0.279 0.455 0.135 0.226 20.146 11.466 4.942 0.444 5.174 0.221 1.596 15.975 11.025 2.522 0.41 4.187 0.822 0.524 18.751 11.502 3.095 0.261 2.791 0.365 0.127 40.657 37.719 17.34 0.286 6.412 0.034 1.343 40.4 28.914 26.566
Australian Elec
MDD ACD SD KD ED DTW SHAP-RE 0.293 1.68 0.088 0.199 19.85 11.8 4.11 0.287 0.887 0.062 0.664 19.864 12.582 4.468 0.28 1.735 0.666 1.679 20.51 10.921 11.67 0.458 4.565 0.136 2.199 15.88 10.933 2.523 0.337 2.658 0.173 0.691 19.327 12.972 4.131 0.277 3.209 0.057 0.527 44.238 40.349 39.504 0.302 6.726 0.365 0.706 41.623 28.13 54.245
ETTH1
MDD ACD SD KD ED DTW SHAP-RE 0.326 1.249 0.355 0.745 19.324 9.475 9.878 0.32 1.376 0.45 0.845 20.103 9.579 13.289 0.29 3.618 0.956 1.5 21.804 10.69 28.806 0.476 2.678 0.371 2.067 16.021 11.35 1.94 0.381 4.233 0.153 0.869 17.323 10.616 7.684 0.291 4.542 0.142 1.0 45.316 38.346 129.161 0.326 4.778 0.751 0.276 25.175 9.745 48.028
ETTH1
MDD ACD SD KD ED DTW SHAP-RE 0.235 2.276 0.104 0.892 20.481 12.525 3.368 0.24 1.163 0.179 0.097 18.292 12.195 3.444 0.225 1.923 1.006 0.651 22.676 12.614 15.306 0.354 4.969 0.332 1.089 15.583 10.499 2.174 0.268 2.67 0.76 0.936 20.463 11.718 5.613 0.254 3.337 0.176 0.645 51.747 49.511 52.871 0.272 6.463 0.134 1.885 42.246 25.674 52.186
Table D.9: Univariate evaluation: similarity metrics reported for Transport datasets.
SDForger Models VAE Models GAN Models Others
ICA3
FPC3
TimeVAE
TimeVQVAE
RTSGAN
SDEGAN
LS4
Bikesharing
MDD ACD SD KD ED DTW SHAP-RE 0.326 0.856 0.101 0.571 21.389 10.305 13.665 0.336 0.691 0.12 0.258 19.406 10.036 13.77 0.295 2.493 0.202 0.125 21.314 10.163 38.651 0.492 1.712 0.042 1.447 16.172 12.804 2.012 0.444 2.451 0.251 0.091 19.806 10.225 19.352 0.29 4.306 0.61 0.218 44.217 31.743 247.88 0.332 1.972 0.035 0.387 21.848 10.347 64.166
Bikesharing
MDD ACD SD KD ED DTW SHAP-RE 0.318 1.653 0.116 0.521 19.761 10.699 4.249 0.322 1.94 0.113 0.806 18.098 10.847 4.63 0.301 4.68 0.569 0.671 25.8 13.416 43.651 0.432 2.715 0.076 1.831 16.097 10.907 2.154 0.365 3.323 0.367 0.511 18.478 10.25 7.005 0.266 4.948 0.425 0.404 44.641 36.872 88.647 0.289 3.845 0.617 1.13 51.221 26.846 100.497
Bikesharing
MDD ACD SD KD ED DTW SHAP-RE 0.37 1.426 0.13 1.91 17.172 12.107 2.807 0.372 1.662 0.488 0.931 20.38 12.47 4.426 0.363 0.086 0.43 0.536 19.349 11.074 7.514 0.451 5.766 0.534 1.776 15.702 10.569 2.189 0.386 2.137 0.808 1.228 20.591 15.251 6.538 0.291 1.244 0.11 0.409 49.379 46.37 42.472 0.339 7.736 0.53 1.334 39.062 26.093 50.348
Traffic
(Junction 1)
MDD ACD SD KD ED DTW SHAP-RE 0.3 0.924 0.276 3.793 18.191 10.776 5.198 0.294 0.861 0.018 2.151 18.465 10.597 5.21 0.277 4.631 0.667 0.769 25.655 13.931 31.453 0.42 2.642 0.008 1.807 15.782 10.893 2.124 0.403 2.195 0.225 0.317 16.151 10.065 4.393 0.267 6.292 0.167 0.052 45.765 41.767 82.871 0.313 3.854 0.266 1.168 36.403 18.485 58.632
Traffic
(Junction 2)
MDD ACD SD KD ED DTW SHAP-RE 0.367 2.893 0.272 1.212 17.937 10.246 5.546 0.368 2.265 0.237 0.645 18.765 10.425 6.777 0.352 2.889 0.526 1.135 21.622 10.307 30.016 0.434 2.878 0.086 1.921 16.007 10.995 2.157 0.445 2.516 0.992 2.667 17.114 11.314 3.426 0.281 5.164 0.328 0.273 44.67 38.389 135.497 0.315 4.492 0.295 1.155 41.478 19.57 99.919
Traffic
(Junction 3)
MDD ACD SD KD ED DTW SHAP-RE 0.326 2.788 0.898 1.086 17.961 11.558 6.848 0.326 2.444 1.039 0.684 18.412 11.147 7.446 0.326 3.861 0.054 0.262 22.054 10.892 46.675 0.41 2.123 0.906 0.336 16.058 11.807 2.022 0.361 3.308 0.23 0.776 18.829 10.677 12.436 0.25 5.97 1.056 1.283 38.533 32.538 118.245 0.276 4.377 1.257 2.822 53.251 25.618 178.736
Table D.10: Univariate evaluation: similarity metrics reported for Nature datasets.
SDForger Models VAE Models GAN Models Others
ICA3
FPC3
TimeVAE
TimeVQVAE
RTSGAN
SDEGAN
LS4
CCP
MDD ACD SD KD ED DTW SHAP-RE 0.173 1.391 0.472 1.644 18.442 15.245 1.123 0.186 1.119 2.21 3.443 18.729 14.344 1.684 0.166 2.203 2.481 4.155 26.232 18.308 13.719 0.3 5.536 1.984 3.31 14.19 10.176 2.486 0.179 2.598 1.12 3.72 14.734 10.831 1.329 0.232 3.06 1.809 5.09 72.136 71.196 52.729 0.236 6.866 1.86 6.164 36.803 28.544 25.99
CCP
MDD ACD SD KD ED DTW SHAP-RE 0.452 0.312 0.416 1.267 16.965 15.495 0.423 0.477 2.223 0.766 0.554 19.28 17.18 1.064 0.409 0.55 0.272 0.212 17.18 14.483 1.13 0.67 6.779 0.441 0.292 15.465 12.381 2.418 0.392 0.56 0.23 0.671 22.525 19.973 2.904 0.364 0.822 0.346 0.011 31.041 30.589 2.805 0.455 9.168 0.753 2.743 35.404 29.251 9.832
CCP
MDD ACD SD KD ED DTW SHAP-RE 0.197 0.662 1.747 2.054 15.92 12.168 1.255 0.191 0.822 2.232 3.638 16.128 11.741 2.426 0.158 0.152 0.007 3.105 15.448 10.881 2.801 0.266 6.672 2.294 4.998 11.295 8.252 2.363 0.189 0.371 0.626 0.251 14.54 10.213 2.042 0.189 0.231 2.729 6.454 38.679 37.706 14.476 0.229 8.524 2.511 8.179 49.348 40.586 47.176
CCP
MDD ACD SD KD ED DTW SHAP-RE 0.178 1.514 1.675 0.157 15.698 11.646 2.299 0.18 1.273 1.638 2.451 17.378 12.004 4.547 0.162 0.535 0.512 0.824 19.159 12.098 6.142 0.249 4.522 1.456 3.091 13.879 10.002 2.228 0.221 2.445 1.398 4.551 21.6 14.647 5.582 0.19 3.142 1.831 4.592 46.177 44.366 31.911 0.249 5.538 1.659 6.204 45.823 33.614 59.928
Table D.11: Univariate evaluation: similarity metrics reported for Finance datasets.
SDForger Models VAE Models GAN Models Others
ICA3
FPC3
TimeVAE
TimeVQVAE
RTSGAN
SDEGAN
LS4
Exchange Rate
(Currency 1)
MDD ACD SD KD ED DTW SHAP-RE 0.27 0.851 0.409 0.32 18.913 15.672 1.759 0.269 1.727 0.62 1.173 19.909 15.91 2.228 0.267 0.211 0.33 0.058 20.652 16.319 2.52 0.415 7.212 0.614 1.77 15.335 11.256 2.449 0.358 1.062 1.031 0.854 26.589 22.394 5.691 0.278 0.932 0.476 0.034 47.853 46.856 17.248 0.308 8.803 0.332 1.275 35.288 26.695 22.698
Exchange Rate
(Currency 2)
MDD ACD SD KD ED DTW SHAP-RE 0.316 0.426 1.285 2.003 17.727 15.12 0.967 0.303 1.424 0.914 1.249 19.579 16.076 1.186 0.286 0.305 0.289 1.507 18.253 14.15 1.744 0.459 7.14 0.863 1.576 14.43 10.689 2.442 0.343 0.983 0.424 0.443 18.272 14.488 2.217 0.253 0.976 1.034 0.208 32.215 31.633 6.681 0.291 8.597 1.083 1.636 45.617 38.042 28.819
Exchange Rate
(Currency 3)
MDD ACD SD KD ED DTW SHAP-RE 0.369 0.113 0.22 0.64 19.578 19.092 0.485 0.428 2.852 0.751 3.039 21.042 19.25 0.956 0.36 0.157 0.122 0.205 19.31 18.532 0.808 0.614 8.035 0.453 2.755 14.765 12.47 2.44 0.47 0.639 1.154 0.046 21.985 20.528 1.168 0.314 0.106 0.633 0.992 28.394 28.143 2.133 0.348 9.934 0.71 0.369 38.289 34.966 15.265
Exchange Rate
(Currency 4)
MDD ACD SD KD ED DTW SHAP-RE 0.342 0.34 0.647 0.484 20.09 16.021 1.18 0.341 1.427 0.252 1.206 20.287 16.187 1.383 0.341 0.848 0.136 0.69 19.622 13.507 1.85 0.543 7.171 0.349 2.483 15.38 10.985 2.34 0.349 1.132 0.49 0.394 20.51 14.745 1.858 0.279 1.023 0.502 0.666 33.675 32.558 7.303 0.317 8.913 0.686 0.645 42.123 33.075 23.273
Exchange Rate
(Currency 6)
MDD ACD SD KD ED DTW SHAP-RE 0.273 0.204 0.285 0.534 19.809 18.469 0.569 0.28 1.893 0.755 3.95 19.763 17.954 0.82 0.251 0.179 0.245 0.88 16.956 14.799 1.036 0.427 7.773 0.342 1.701 14.579 11.822 2.386 0.338 0.578 0.049 1.127 21.034 18.02 2.427 0.237 0.414 0.481 0.152 39.025 38.671 5.158 0.286 8.341 0.464 1.401 34.643 30.317 28.882
Exchange Rate
(Currency 7)
MDD ACD SD KD ED DTW SHAP-RE 0.318 0.17 0.249 0.743 19.565 18.484 0.945 0.353 2.423 0.169 3.389 23.189 20.773 0.761 0.322 0.179 0.212 0.574 15.034 13.105 0.981 0.493 7.967 0.22 2.616 14.768 12.875 2.496 0.374 0.121 0.61 0.379 21.199 19.659 0.918 0.292 0.22 0.344 0.759 36.463 36.214 2.649 0.334 9.778 0.599 0.443 34.437 30.555 9.963
Exchange Rate
(Currency 8)
MDD ACD SD KD ED DTW SHAP-RE 0.401 0.06 0.091 1.009 16.88 16.254 0.232 0.378 2.259 0.04 2.867 21.571 20.245 1.044 0.358 0.119 0.595 0.716 16.902 15.651 0.638 0.564 8.052 0.173 2.751 14.882 12.884 2.43 0.467 0.197 0.253 0.202 21.432 20.061 0.679 0.339 0.077 0.041 0.891 37.487 37.232 3.228 0.353 10.102 0.333 0.205 42.048 39.402 15.895

### v2

@@
 By enabling textual conditioning in the generation process, SDForger paves the way for multimodal modeling and the streamlined integration of time series with textual information.
-SDForger source code will be open-sourced soon.
+The model is open-sourced at https://github.com/IBM/fms-dgt/tree/main/fms_dgt/public/databuilders/time_series.
 1 Introduction
@@
 Despite these innovations, most existing approaches either require task-specific pretraining or struggle to model complex structures and may benefit from synthetic data for improving generalization and robustness.
+Specialized architectures for time-series generation
+Several architectures have been specifically designed to capture temporal and multivariate dependencies in time series.
+Variational autoencoders such as TimeVAE (Desai et al., 2021) and TimeVQVAE (Lee et al., 2023) use recurrent or vector-quantized structures to model sequential dynamics.
+GAN-based approaches, including RTSGAN (Pei et al., 2021), SDEGAN (Kidger et al., 2021), and COSCI-GAN (Seyfi et al., 2022), employ recurrent or component-wise disentangled generators to capture complex temporal patterns adversarially.
+Diffusion-based models, like LS4 (Zhou et al., 2023), generate sequences through learned reverse-time processes.
+These specialized architectures complement general-purpose time-series generation methods and provide valuable baselines for evaluating synthetic data.
+However, these architectures are trained from scratch and cannot leverage existing pre-trained language or foundation models, limiting their scalability and adaptability across domains.
 3 Methodology
@@
 We denote the $i$-th row of $E$ as $E_{i}$, corresponding to the embedding vector of instance $X_{i}$, and the value in its $k$-th column as $e_{ik}$.
-More details on the choice of $k_{c}$ are given in Appendix A.2.
+More details on the choice of $k_{c}$ are given in Appendix A.2 and Appendix Table D.1.
 Notably, both methods offer the advantage that their computational cost depends on the number of instances $I$ and the number of components $k_{c}$, but not on the instance length $L$.
@@
 Let $\mathcal{P}^{\textup{FT}}=\{{\mathcal{P}_{i}^{\textup{FT}}\}}_{i=1}^{I}$ denote the set of fine-tuning prompts, where:
-[…]
-[…]
+$\mathcal{P}_{i}^{\textup{FT}}=\texttt{``}$ $\texttt{Input:\ }\bigcirc_{k=1}^{K}{\texttt{(value\_{$\pi(k)$} is [blank],)}}\texttt{\ [sep]\ }$
+$\texttt{Target:\ }\bigcirc_{k=1}^{K}{\texttt{($e_{i\pi(k)}$ [answer])''}}$
 […]
@@
 Given the embedding table $E\in\mathbb{R}^{I\times K}$, we define the set of inference prompts at each generation step as $\mathcal{P}^{\textup{INF}}=\{{\mathcal{P}_{g}^{\textup{INF}}\}}_{g=1}^{G}$ where:
-[…]
+$\mathcal{P}_{g}^{\textup{INF}}=\texttt{``}$ $\texttt{Input:}\bigcirc_{k=1}^{K}{\texttt{( value\_{$\pi(k)$} is [blank], ) }}\texttt{ [sep] }\texttt{Target:'' }$
 We use a multinomial distribution sampling strategy to reduce repetition and generate more creative and diverse outputs.
@@
 At each inference step, we generate a batch of $G$ synthetic instances, repeating the process until the desired number of sequences is obtained or a stopping criterion is met.
-We denote the set of all generated text instances as: $\mathcal{G}=\{\mathcal{G}_{1},\dots,\mathcal{G}_{G}\}.$ Ideally, the fine-tuned LLM should generate text instances in the following format: $\mathcal{G}_{g}=\bigcirc\ (\ \mathcal{P}_{g}^{\textup{INF}},\ \bigcirc_{k=1}^{K}(\text{$\tilde{a}_{g\pi(k)}$ {{[answer]}}}\ )\ )$] ) ) where $\pi$ is the random permutation used in $\mathcal{P}_{g}^{\textup{INF}}$, and $\{\tilde{a}_{g\pi(k)}\}_{k=1}^{K}$ are the $K$ numerical values inferred, which form the generated embedding table.
+We denote the set of all generated text instances as: $\mathcal{G}=\{\mathcal{G}_{1},\dots,\mathcal{G}_{G}\}.$ Ideally, the fine-tuned LLM should generate text instances in the following format: $\mathcal{G}_{g}=\bigcirc\ (\ \mathcal{P}_{g}^{\textup{INF}},\ \bigcirc_{k=1}^{K}(\text{$\tilde{a}_{g\pi(k)}$ {{[answer]}}}\ )\ )$ where $\pi$ is the random permutation used in $\mathcal{P}_{g}^{\textup{INF}}$, and $\{\tilde{a}_{g\pi(k)}\}_{k=1}^{K}$ are the $K$ numerical values inferred, which form the generated embedding table.
 Retrieve embedding from text
@@
 At each inference step $s$, we apply an online filtering procedure that validates generated embeddings without requiring reconstruction into the time-series domain, ensuring efficient real-time evaluation.
-Specifically, the reconstructed embedding matrices $\big{(}\tilde{E}^{1,s},\dots,\tilde{E}^{C,s}\big{)}$ are filtered based on three criteria: 1) Instances with missing values are discarded, as they prevent accurate reconstruction; 2) Duplicated instances are discarded to maintain diversity in the generated dataset; 3) Significantly diverging instances are discarded (procedure in Appendix A.3).
+Specifically, the reconstructed embedding matrices $\big(\tilde{E}^{1,s},\dots,\tilde{E}^{C,s}\big)$ are filtered based on three criteria: 1) Instances with missing values are discarded, as they prevent accurate reconstruction; 2) Duplicated instances are discarded to maintain diversity in the generated dataset; 3) Significantly diverging instances are discarded.
 This combined filtering procedure not only enforces diversity and validity among the generated instances but also provides a diagnostic signal: if a substantial fraction of samples is rejected, it may indicate that the fine-tuned LLM requires further training or more representative data.
+Representative examples of discarded instances and details on the divergence detection procedure are provided in Appendix A.3, while Appendix Table D.2 reports the rejection rates observed in a representative generation scenario, illustrating the balance between filtering rigor and sample diversity.
 In generation mode, SDForger employs a dynamic stopping criterion that continues generating batches of $G$ text instances as long as sufficient diversity is preserved among the generated samples (Appendix A.4).
@@
 We evaluated SDForger models using 12 publicly available datasets from various domains, including energy, transport, industry, weather, and finance, with sampling frequencies ranging from 2 minutes to monthly.
-The datasets were sourced from the Monash Time Series Forecasting Repository and public domains.
+The datasets, sourced from the Monash Time Series Forecasting Repository and other public domains, include both stationary and non-stationary time series, reflecting diverse temporal dynamics.
 Detailed information is provided in Appendix C.
@@
 5.1 Similarity-based metrics results
-The similarity-based results aggregated for the multisample and univariate settings are reported in Table 1, with detailed per-dataset scores provided in Appendix Tables D.7, D.8, D.9, D.10, and D.11.
-Overall performance While individual baselines may excel on specific metrics or datasets, SDForger stands out for its consistently strong and balanced performance across all metrics and domains.
-This pattern is particularly notable for the ICA-based variant, which avoids the common trade-off between statistical and structural fidelity.
-SDForger’s ability to maintain competitive scores across diverse metrics, as further supported by the average rank scores, underscores its generalization and robustness across heterogeneous temporal datasets.
-By decoupling representation learning from generation, SDForger is able to capture long-range dependencies while maintaining fidelity to statistical features.
-This design appears especially effective in producing temporally coherent, domain-consistent samples.
+The similarity-based results aggregated for the multisample and univariate settings are reported in Table 1, with detailed per-dataset scores provided in Appendix Tables D.10, D.11, D.12, D.13, and D.14.
+Overall performance.
+Different generative models exhibit complementary strengths: for instance, TimeVAE performs well on distribution-based metrics, while TimeVQVAE excels on distance-based measures such as Euclidean Distance and DTW.
+In contrast, SDForger achieves consistently strong and balanced performance across both metric categories, maintaining high scores without overfitting to either statistical or structural similarity (Table 1).
+This balanced behaviour is further confirmed by the normalized average scores per metric group and the average rank values.
+Such consistency indicates that SDForger not only preserves key statistical features but also captures the underlying temporal and distributional structure of the data, demonstrating strong generalization and robustness across heterogeneous temporal domains.
+By decoupling representation learning from generation, SDForger captures long-range dependencies while maintaining statistical realism, ultimately producing temporally coherent and domain-consistent synthetic samples.
 Robustness to evaluation protocols Comparing multisample and univariate settings, we observe that model rankings and relative performances remain largely consistent, suggesting that SDForger is robust to variations in the evaluation protocol.
@@
 ICA vs.
-FPC Between the two embedding strategies, ICA consistently leads, particularly on distance-based metrics.
-This may reflect the greater interpretability and statistical independence of ICA components, which offer more stable conditioning inputs for generation.
+FPC The ICA embedding strategy consistently leads, particularly on distance-based metrics.
+The superior performance of the ICA-based variant likely comes from the nature of the components it produces.
+Unlike FPC, which orders components by explained variance and often concentrates most information in the first few components, ICA explicitly seeks statistically independent components.
+This tends to produce a more balanced and disentangled basis decomposition, where each component carries distinct information that have similar importance for data reconstruction.
+For our LLM-based generation pipeline, this disentanglement appears advantageous because the model can learn a joint distribution over a set of factors that all have the same “power”.
+Thus, the results suggest that the LLM is indeed better equipped to model the joint distribution when presented with independent factors rather than a hierarchy of variance-ordered components.
 Table 1: Aggregated performance comparison in the multisample and univariate settings.
-Metrics include raw similarity scores and normalized averages (in $[0-1]$) for each metric group, plus the average rank.
+Metrics include raw similarity scores and normalized averages (in [01][0-1]) for each metric group, plus the average rank.
 Lower values are better.
@@
 For LS4, we modified the architecture to support multivariate generation.
-We also fine-tune the embedding dimension $k$ in SDForger based on dataset complexity: $k{=}3$ for bikesharing, $k{=}7$ for etth1, and $k{=}5$ for traffic.
+Furthermore, we adjust the embedding dimension $k$ in SDForger according to the complexity of each dataset, setting k=3k{=}3 for bikesharing, k=7k{=}7 for etth1, and k=5k{=}5 for traffic.
+To determine these values, we conducted a small ablation study to identify the optimal embedding dimension for each dataset (see Table D.7).
 SDForger demonstrates strong performance across datasets, as evidenced by its top average rank, with notable results on bikesharing and etth1.
@@
 Effect of embedding dimension $\bm{k}$ Appendix Table D.6 presents an ablation study on the number of components $k$ used in SDForger’s embedding space.
-A compact embedding with $k{=}3$ offers strong performance across both multisample and univariate settings, indicating that a small number of components is often sufficient to capture core temporal and structural patterns.
+A compact embedding with k=3k{=}3 offers strong performance across both multisample and univariate settings, indicating that a small number of components is often sufficient to capture core temporal and structural patterns.
 However, the optimal value of $k$ may vary across datasets, with more intricate dynamics potentially requiring higher-dimensional representations.
 In practice, users may also opt to select $k$ based on a desired percentage of explained variance, adapting the representation to specific application needs.
-Domain-level insights Appendix Tables D.1 and D.2 summarize the average normalized similarity scores per dataset.
+Domain-level insights Appendix Tables D.3 and D.4 summarize the average normalized similarity scores per dataset.
 SDForger models achieve strong performance across a wide range of domains.
@@
 Nonetheless, SDForger maintains competitive results even in more complex settings, underscoring the flexibility of the proposed architecture.
-Generation efficiency Appendix Table D.3 reports the average time to generate univariate sequences for three targets from the Bikesharing dataset, across two window lengths.
-SDForger is substantially faster than all competitors, often by one to two orders of magnitude.
-TimeVAE is the closest competitor but remains over 4$\times$ slower.
-Notably, unlike GAN-based competitors, SDForger’s generation time is independent of sequence length and scales with the number of embedding components ($k$).
-A minor exception occurs at $k=3$, where the reduced latent expressivity increases LLM fine-tuning time.
-Overall, SDForger achieves state-of-the-art efficiency without compromising quality.
+[…]
 LLM comparison GPT-2 (124M) achieves performance on par with, and sometimes better than larger and more recent models such as granite-3.0 (2B) and phi-3.5 (3.8B) (Appendix Table D.9).
@@
 While runtime cost grows with model size, SDForger remains efficient compared to baselines (Appendix Table D.8).
+Filtering Procedure.
+Appendix Table D.2 reports rejection statistics for a representative generation scenario.
+We observe that the overall discard rate remains consistently low ($2\%$) across settings, indicating that most generated embeddings fall within a plausible norm range.
+The proportion of missing values increases with larger embedding dimensions, reflecting a higher likelihood of incomplete generations in longer textual outputs.
+Notably, the $\ell_{2}$ norms of accepted samples closely match those of the original embeddings, while discarded ones exhibit markedly higher values, confirming that the filtering procedure effectively removes divergent or anomalous generations.
 6 Shaping time series with language
@@
 Moreover, by leveraging LLMs, SDForger enables seamless integration with textual prompts, paving the way for multimodal time-series generation, where natural language can guide not only content but also structure, semantics, or temporal context.
-One limitation of our approach is the trade-off between embedding dimensionality and training set size: using too many components relative to the number of instances can lead to unstable fine-tuning and reduced generation quality.
-Additionally, our current implementation supports only autoregressive LLMs; future work could extend the framework to encoder-only models and different generation paradigms such as masked token prediction.
-Further improvements may also come from adaptive embedding strategies and adding contextual information (Section 6) for more reliable generation.
+We believe SDForger can be further improved.
+Its modular design is intentionally built to support flexible experimentation, making it easy to explore enhancements or tailor components to specific needs.
+We see several promising directions:
+  • Embedding Strategies While our current approach relies on linear methods like FastICA and FPC, future work could explore more expressive, nonlinear embeddings (e.g., AE) or multivariate-aware methods like Multivariate FPCA or Multivariate Singular Spectrum Analysis to better capture temporal and inter-channel dependencies.
+  • Parameter-Efficient Fine-Tuning We currently use full fine-tuning for the LLM.
+However, using too many components relative to the number of instances can lead to unstable fine-tuning and reduced generation quality.
+Incorporating PEFT techniques such as LoRA or adapters could improve scalability, efficiency, and facilitate domain adaptation.
+  • Extension to encoder-only models Our current implementation supports only autoregressive LLMs; future work would extend the framework to encoder-only models and different generation paradigms such as masked token prediction.
+  • Extended Utility Evaluation While we focus on forecasting, SDForger could be evaluated and optimized for broader downstream tasks such as classification or anomaly detection.
+  • Context and Covariate Integration By design, SDForger supports integration of external covariates (e.g., categorical or textual data).
+Expanding this functionality could enable richer conditional generation, and multimodal transfer learning (see Section 6).
+In summary, SDForger offers a flexible foundation, and we see meaningful opportunities to improve it both architecturally and in terms of task generalization.
 Appendix
@@
   1. 1.
-    Compute the ACF and identify significant peaks, excluding lag 00.
+    Compute the ACF and identify significant peaks, excluding lag 0.
   2. 2.
@@
 However, this property also makes FastICA embeddings more robust during generation, as information is distributed more evenly across components, reducing the risk that a few badly generated components disproportionately affect the synthesized curves.
+To provide a quantitative intuition, Appendix Table D.1 reports the proportion of variance retained across embedding dimensions for both decomposition methods.
+This analysis highlights how the variance explained increases with $k$, and how FPC typically achieves higher cumulative variance with fewer components, while FastICA distributes information more evenly across dimensions.
 In practice, we recommend keeping the total number of components $K$ reasonably small, particularly when the number of training instances is limited.
 Empirically, with a training set of 30 instances, setting $K>25$ often results in unstable fine-tuning and an increased rate of discarded samples due to low-quality generation.
-This limitation stems from the LLM’s reduced ability to effectively model high-dimensional embeddings under data-scarce conditions.
+This limitation stems from the LLM’s reduced ability to model high-dimensional embeddings under data-scarce conditions effectively.
 A.3 In-generation filtering
+Missing values and duplicated instances.
+To illustrate the filtering logic, we report three concrete examples of generated prompts with embedding dimension $K=4$. Following the inference template:
+$\mathcal{P}_{g}^{\textup{INF}}=\texttt{``Input:}\bigcirc_{k=1}^{K}\texttt{( value\_{$\pi(k)$} is [blank], )}\texttt{ [sep] Target:''},$
+the model produces the following textual generations:
+  • Prompt 1 (valid)
+    Input: value_2 is [blank], value_4 is [blank], value_1 is [blank], value_3 is [blank] [sep] Target: 0.125 [answer] -0.084 [answer] 0.217 [answer] 0.041 [answer]
+  • Prompt 2 (duplicated)
+    Input: value_4 is [blank], value_1 is [blank], value_2 is [blank], value_3 is [blank] [sep] Target: -0.084 [answer] 0.217 [answer] 0.125 [answer] 0.041 [answer]
+  • Prompt 3 (missing value)
+    Input: value_1 is [blank], value_3 is [blank], value_2 is [blank], value_4 is [blank] [sep] Target: 0.182 [answer] 0.095 [answer] -0.012 [answer]
+In this example, the filtering stage identifies Prompt 2 as a duplicate of Prompt 1 and discards it, while Prompt 3 is removed because it does not contain all the targets’ coefficients.
+Only Prompt 1 is retained for reconstruction.
+This simple yet effective procedure ensures that the generated embedding tables remain diverse and valid before decoding into the time-series domain.
+Diverging instances
 To ensure the quality of synthetic data, we discard generated instances whose embedding coefficients significantly deviate from the distribution of the original data.
@@
 This criterion efficiently filters out extreme outliers in the latent space, without requiring reconstruction into the time-series domain.
-Formally, for each channel $c$, let $\hat{E}^{c,s}=E^{c}\cup\tilde{E}^{c,\leq s-1}$ be the matrix containing both original embeddings and all previously accepted generated embeddings up to inference step $s-1$, with $\hat{E}^{c,0}=E^{c}$.
+Formally, for each channel $c$, let $\hat{E}^{c,s}=E^{c}\cup\tilde{E}^{c,\leq s-1}$ be the matrix containing both original embeddings and all previously accepted generated embeddings up to inference step s1s-1, with $\hat{E}^{c,0}=E^{c}$.
 Denote by $N^{c,\text{old}}$ and $N^{c,\text{new}}$ the sets of squared Euclidean norms of the rows of $\hat{E}^{c,s-1}$ and the newly generated matrix $\tilde{E}^{c,s}$, respectively.
@@
 B.1 Feature-based evaluation
-Marginal Distribution Difference.
+Marginal Distribution Difference
 MDD computes an empirical histogram for each dimension and time step in the generated series, using the bin centers and widths from the original series.
 It then calculates the average absolute difference between this histogram and that of the original series across bins, assessing how closely the distributions of the original and generated series align.
-AutoCorrelation Difference.
+AutoCorrelation Difference
 ACD computes the autocorrelation of both the original and generated time series, then determines their difference.
 By contrasting the autocorrelations, we could evaluate how well dependencies are maintained in the generated time series.
-Skewness Difference.
+Skewness Difference
 SD is vital for the marginal distribution of a time series, quantifying its distribution asymmetry.
@@
 $SD=\left|\frac{\mathbb{E}[(T_{s}^{gen}-\mu_{s}^{gen})^{3}]}{\sigma_{s}^{gen3}}-\frac{\mathbb{E}[(T_{s}^{tr}-\mu_{s}^{tr})^{3}]}{\sigma_{s}^{tr3}}\right|.$
-Kurtosis Difference.
+Kurtosis Difference
 Like skewness, KD assesses the tail behavior of a distribution, revealing extreme deviations from the mean.
@@
 B.2 Distance-based evaluation
-Euclidean Distance.
+Euclidean Distance
 For each original series $s^{tr}=(x_{1},...,x_{l})$ and its generated $s^{gen}=(y_{1},...,y_{l})$, $ED=\sqrt{\sum_{1=1}^{l}(x_{i}-y_{i})^{2}}$.
@@
 It provides a value-wise comparison between the time series.
-Dynamic Time Warping.
+Dynamic Time Warping
 Given that ED overlooks alignment, we include DTW to capture the optimal alignment between series regardless of their pace or timing.
 The alignment facilitated by DTW offers insights into the predictive quality of the generated series.
-Shapelet-based Reconstructions.
+Shapelet-based Reconstructions
 Shapelet based RE is calculated by generated time series using shapelets extracted using shift invariant dictionary learning (SIDL) algorithm (Zheng et al., 2016).
@@
   • Exchange Rate (Godahewa et al., 2021) contains daily exchange rates for currencies of eight countries (Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore) between 1990 and 2016.
-  • NN5 (Daily, Weekly) (Godahewa et al., 2021) contains cash withdrawal data from ATMs.
+This dataset is non-stationary.
+  • NN5 (Daily, Weekly) (Godahewa et al., 2021) contains cash withdrawal data from ATMs. This dataset combines stationary and non-stationary time series.
 C.2 Protocols
@@
 Appendix D Additional results
+Table D.1: Variance retained across embedding dimensions.
+For each dataset, we report the proportion of total variance retained for embedding dimensions $k=3$, $k=5$, and $k=7$ under both FastICA and FPC decompositions.
+Dataset FICA FPC
+k=3 k=5 k=7 k=3 k=5 k=7
+Appliances 0.333 0.457 0.539 0.324 0.459 0.555
+Australian Electricity 0.703 0.853 0.912 0.787 0.895 0.938
+Bikesharing 0.492 0.662 0.744 0.599 0.734 0.796
+Carbon Capture Plant 0.773 0.905 0.951 0.843 0.933 0.965
+ETTH1 0.536 0.686 0.775 0.635 0.754 0.823
+ECL 0.809 0.941 0.971 0.991 0.997 0.999
+Exchange Rate 0.776 0.907 0.942 0.957 0.982 0.989
+NN5 0.336 0.479 0.590 0.719 0.780 0.826
+Tourism 0.869 0.954 0.977 0.986 0.995 0.998
+Traffic 0.376 0.564 0.697 0.408 0.591 0.714
+Traffic Monash 0.541 0.722 0.823 0.745 0.845 0.902
+Rain - Weather 0.449 0.605 0.725 0.558 0.684 0.781
+Solar - Weather 0.450 0.616 0.718 0.674 0.773 0.833
+Temperature Max - Weather 0.493 0.602 0.685 0.864 0.893 0.915
+Temperature Min - Weather 0.390 0.517 0.611 0.777 0.824 0.858
+Table D.2: Filtering statistics for generated embeddings.
+Rejection statistics on the count variable from the bikesharing dataset, averaged across 5 seeds.
+Each row reports the proportion of generated samples containing missing values, the fraction of samples discarded by the filtering stage, and the average $\ell_{2}$ norms of the original, accepted, and discarded embedding vectors.
+NaN% Discard% Norms Original (Avg) Norms Accepted (Avg) Norms Discarded (Avg)
+SDF-ICA3 3.87 1.94 1.708 1.714 19.066
+SDF-ICA5 36.29 1.94 2.202 2.248 9.514
+SDF-ICA7 58.82 0.00 2.602 2.521 0.000
 Table D.3: Per-dataset similarity results in the multisample setting. Average normalized similarity scores (feature-based and distance-based) for each dataset and model.
@@
 SDF-ICA7 0.306 0.597 0.736 1.458 19.989 12.381 8.102 0.175 0.543
+Table D.7: Ablation study: embedding dimension.
+TTM forecasting performance on downstream tasks using different training sources: generated data, and a combination of original and generated data.
+Results are reported for 3 multivariate datasets: bikesharing (target: count, control: temperature, humidity), etth1 (target: HUFL, control: MUFL, OT), and traffic (target: junction1, control: junction2, junction3).
+Metrics include RMSE, MASE, WQL, and average rank (lower is better).
+Bold highlights the best result within each row group; bold+underlined the overall best.
+bikesharing etth1 traffic
+RMSE MASE WQL RMSE MASE WQL RMSE MASE WQL
+0-shot 0.728 2.150 0.287 0.678 2.132 0.255 0.708 1.555 0.255
+Original Data (OD) 0.495 0.822 0.178 0.658 1.820 0.232 0.702 1.995 0.283
+GEN SDF-FPC3 0.527 0.926 0.200 0.692 1.914 0.246 0.699 2.029 0.287
+SDF-FPC5 0.530 0.918 0.198 0.693 2.003 0.252 0.662 1.837 0.262
+SDF-FPC7 0.522 0.915 0.197 0.650 1.887 0.232 0.812 2.265 0.323
+SDF-ICA3 0.514 0.899 0.194 0.647 1.829 0.233 0.730 2.068 0.294
+SDF-ICA5 0.537 0.909 0.194 0.637 1.934 0.233 0.655 1.849 0.262
+SDF-ICA7 0.517 0.898 0.193 0.626 1.820 0.224 0.790 2.189 0.312
+OG + GEN SDF-FPC3 + OD 0.493 0.829 0.179 0.658 1.780 0.229 0.736 2.077 0.296
+SDF-FPC5 + OD 0.487 0.807 0.174 0.659 1.757 0.230 0.743 2.087 0.297
+SDF-FPC7 + OD 0.492 0.821 0.177 0.666 1.754 0.231 0.706 1.993 0.283
+SDF-ICA3 + OD 0.487 0.801 0.173 0.640 1.790 0.228 0.734 2.074 0.295
+SDF-ICA5 + OD 0.486 0.804 0.174 0.649 1.780 0.230 0.750 2.110 0.301
+SDF-ICA7 + OD 0.490 0.810 0.175 0.642 1.746 0.226 0.718 2.025 0.288
 Table D.8: Average Generation Time Across LLM Backbones.